OpenStudy (anonymous):

Can somebody explain to me why cos2x = cos2x - sin2x?

OpenStudy (bahrom7893):

You mean Cos(2x) = Cos^2(x) - Sin^2(x)

OpenStudy (bahrom7893):

Look up the formula for Cos(A+B) and post it here.

OpenStudy (anonymous):

Oh, ok. cos(A+B) = cos(A)cos(B) - sin(A)sin(B)

OpenStudy (anonymous):

Oh, yeah. Sorry. It is squared. Cos(2x) = Cos^2(x) - Sin^2(x)

OpenStudy (bahrom7893):

Okay, so in your case you have: Cos(2x) = Cos(x+x) = Cos(x)Cos(x) - Sin(x)Sin(x) = ...

OpenStudy (bahrom7893):

Do you get it?

OpenStudy (anonymous):

Ah. I see. Thanks.

OpenStudy (bahrom7893):

Np